# Countably compact inverse semigroups and Nyikos problem

### v1

@@ 4. Nyikos semilattices
To derive a contradiction assume that $X$ contains a chain $L=\{l_{\alpha}:\alpha<\omega_{1}\}$ such that $l_{\alpha}\leq l_{\beta}$ if and only if $\alpha\leq\beta$. By [22, Corollary 3.10.15], $X{\times}X$ is countable compact and thus pseudocompact. Theorem 2.9 implies that $\beta X$ is a compact topological semigroup. Since $X$ is dense in $\beta X$, we obtain that $\beta X$ is a semilattice. By Lemma 3.4, there exists $z=\sup L\in\beta X$. Lemma 4.7 implies that there exists $l_{\alpha}\in L$ such that for each open neighborhood $U$ of $l_{\alpha}$ the set $U\setminus I_{L}$ contains an element $d_{U}$. Since $X$ is locally compact, there exists an open neighborhood $W$ of $l_{\alpha}$ such that $\operatorname{cl}_{X}(W)$ is compact. The continuity of the semilattice operation in $\beta X$ yields open neighborhoods $U$ of $l_{\alpha}$ and $V$ of $z$ such that $UV\subseteq W$. By Lemma 3.4, there exists an ordinal $\delta<\omega_{1}$ such that $L^{\prime}=\{l_{\xi}:\delta<\xi<\omega_{1}\}\subset V$. Note that $d_{U}L^{\prime}\subseteq W$, $L^{\prime}$ is isomorphic to $(\omega_{1},\min)$ and $|d_{U}L^{\prime}|=|d_{U}L|=\omega_{1}$, by the choice of $d_{U}$. Since for each $a\in X$ the map $\phi_{a}:X\rightarrow X$, $\phi_{a}(x)=ax$ is a homomorphism, Lemma 3.7 implies that the chain $d_{U}L^{\prime}$ is isomorphic to $(\beta,\min)$ for some ordinal $\beta\leq\omega_{1}$. Taking into account that $|d_{U}L^{\prime}|=\omega_{1}$, we get that $d_{U}L^{\prime}$ is isomorphic to $(\omega_{1},\min)$. Since $\operatorname{cl}_{X}(W)$ is compact, the chain $\operatorname{cl}_{X}(d_{U}L^{\prime})\subseteq\operatorname{cl}_{X}(W)$ is compact, which contradicts Lemma 3.6. ∎
@@
Seeking a contradiction, assume that a locally compact Nyikos semilattice $X$ contains a chain $L=\{l_{\alpha}:\alpha<\omega_{1}\}$ such that $l_{\alpha}\leq l_{\beta}$ if and only if $\alpha\geq\beta$. By Theorem 2.9, $\beta X$ is a compact topological semigroup. Since $X$ is dense in $\beta X$, we obtain that $\beta X$ is a semilattice. By Lemma 3.5, there exists $z=\inf L\in\beta X\setminus X$.
@@
Lemma 4.5 implies that $I_{L}$ is closed. To derive a contradiction, assume that $I_{L}$ is not open. Then there exists $y\in I_{L}$ such that each open neighborhood $U$ of $y$ contains an element $d_{U}\in X\setminus I_{L}$. Let us show that $yz\in X$. Since $y\in I_{L}$, there exists an ordinal $\alpha<\omega_{1}$ such that $yl_{\xi}=yl_{\alpha}$ for each $\xi\geq\alpha$. By Lemma 3.5, for each open neighborhood $V$ of $z$ there exists $\delta_{V}<\omega_{1}$ such that $\{l_{\xi}:\xi\geq\delta_{V}\}\subset V$. Then $yl_{\alpha}\in yV$ for each open neighborhood $V\subset\beta X$ of $z$. Since $\beta X$ is a Hausdorff topological semilattice, we get $yz=yl_{\alpha}\in X$. Since $X$ is locally compact, there is an open neighborhood $W\subseteq X$ of $yz$ such that $\operatorname{cl}_{X}(W)$ is compact. The continuity of the semilattice operation in $\beta X$ yields open neighborhoods $U$ of $y$ and $V$ of $z$ such that $UV\subseteq W$. Let $L^{\prime}=\{l_{\alpha}:\alpha\geq\delta_{V}\}\subset V$. Note that $d_{U}L^{\prime}\subseteq W$. By the choice of $d_{U}$, $|d_{U}L|=\omega_{1}$, which implies $|d_{U}L^{\prime}|=\omega_{1}$. As left shifts in $\beta X$ are homomorphisms and $L^{\prime}$ is order isomorphic to $(\omega_{1},\max)$, Lemma 3.7(2) implies that $d_{U}L^{\prime}$ is also order isomorphic to $(\omega_{1},\max)$. Since $\operatorname{cl}_{X}(W)$ is compact, the chain $\operatorname{cl}_{X}(d_{U}L^{\prime})\subset\operatorname{cl}_{X}(W)$ is compact, which contradicts Lemma 3.6. ∎
By [22, Corollary 3.6.5], the set $\operatorname{cl}_{\beta X}(I_{L})$ is clopen.
@@
There exists a finite set $F\subset X\setminus I_{L}$ such that $\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})=\bigcup_{x\in F}{\uparrow}x$.
@@
Let us first check that $\operatorname{cl}_{\beta X}(I_{L})$ is an ideal in $\beta X$. It suffices to show that $ab\in\operatorname{cl}_{\beta X}(I_{L})$ for any $a\in\operatorname{cl}_{\beta X}(I_{L})$ and $b\in\beta X$. Fix any open neighborhood $W$ of $ab$. By the continuity of the semigroup operation in $\beta X$, there exist open neighborhoods $U$ of $a$ and $V$ of $b$ such that $UV\subseteq W$. Then there exist $c\in U\cap I_{L}$ and $d\in X\cap V$ such that $cd\in W\cap I_{L}$. Thus $ab\in\operatorname{cl}_{\beta X}(I_{L})$, as required. Consider the Rees quotient semilattice $S=\beta X/{\operatorname{cl}_{\beta X}(I_{L})}$ which is obtained by contracting the clopen ideal $\operatorname{cl}_{\beta X}(I_{L})$ to a point denoted by 00. Obviously, $S$ is a compact topological semilattice, $0=\inf S$ and 00 is isolated in $S$. Further we agree to identify the sets $S\setminus\{0\}$ and $\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})$. Observe that the set $\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})$ is nonempty, as it contains $L$. Let $\mathfrak{C}$ be the set of all maximal chains in $S\setminus\{0\}$. By the compactness of $S\setminus\{0\}$, each $C\in\mathfrak{C}$ contains its minimum. To derive a contradiction, assume that the set
$F=\{\min C:C\in\mathfrak{C}\}\subset\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})$
is infinite. It is easy to see that $ab=0$ for any distinct $a,b\in F$. So, $T=F\cup\{0\}$ is a subsemilattice of $S$. Since each chain in $T$ is finite, Theorem 3.2 implies that $T$ is closed in $S$ and thus compact. Consider an accumulation point $y\in T$ of $F$. Observe that for each open neighborhood $U$ of $y$ there exist distinct $a,b\in U\cap F$ witnessing that $0=ab\in UU$. Since $S$ is a Hausdorff topological semilattice, we get that $y=yy=0$. On the other hand, $y\neq 0$, as 00 is an isolated point, a contradiction. So the set $F$ is finite. At this point it is straightforward to check that $S\setminus\{0\}=\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})=\bigcup_{x\in F}{\uparrow}x$.
It remains to check that $F\subset X$. Fix any $f\in F$. Observe that
@@
is an open subsemilattice of $S$ such that $\inf P=f\in P$. Then $P\cap X$ is an open subsemilattice of $X$. Taking into account that $X$ is separable, Lemma 4.1 yields that the semilattice $\operatorname{cl}_{X}(P\cap X)$ contains the smallest element $p$. Since $\operatorname{cl}_{X}(P\cap X)\subseteq\operatorname{cl}_{\beta X}(P)\subseteq{\uparrow}f$, we get $f\leq p$. Since $X$ is dense in $\beta X$ and $P$ is open, for every open neighborhood $U$ of $f$ the set $U\cap P\cap X$ is not empty. Thus $p\in Up$ for each open neighborhood $U$ of $f$. Since $S$ is a Hausdorff topological semilattice, $fp=p$, i.e. $p\leq f$. Hence $p=f$ and $f\in\operatorname{cl}_{X}(P\cap X)\subset X$. ∎
@@
Since $L\subset X\setminus I_{L}$, we obtain
@@
By Claim 4.11, there exists $f\in F\subset X\setminus I_{L}$ such that $f\leq z$. But then $fL=(fz)L=f(zL)=\{fz\}=\{f\}$. Hence $f\in I_{L}$, a contradiction. ∎
@@
Propositions 4.8 and 4.9 imply that each locally compact Nyikos topological semilattice $X$ contains no isomorphic copies of $(\omega_{1},\min)$ or $(\omega_{1},\max)$. By Theorem B, $X$ is compact. ∎
5. Proof of the main result and final remarks
@@
For more about Green’s relations on inverse semigroups see [36, Chapter 3.2]. The following two results can be considered folklore.
@@
In fact, the proof of Proposition 5.2 follows from the proof of [31, Lemma 2.2.3]. We are in a position to prove Theorem A, which states that every locally compact Nyikos inverse topological semigroup is compact.
Proof of Theorem A.
Consider a locally compact Nyikos inverse topological semigroup $S$. By Corollary 1.4, $S$ is a topological inverse semigroup. Note that the semilattice of idempotents $E(S)$ is a retract of $S$ under the continuous map $\pi:x\mapsto xx^{-1}$. It follows that $E(S)$ is a locally compact Nyikos topological semilattice. Theorem 4.12 implies that $E(S)$ is compact. By Theorem C, $\beta S$ is a topological inverse semigroup. Since $S$ is dense in $\beta S$, we get that $E(S)$ is dense in $E(\beta S)$. Thus $E(S)=E(\beta S)$.
Seeking a contradiction, assume that $\beta S\setminus S\neq\varnothing$. Since $S$ is locally compact, the remainder $\beta S\setminus S$ is closed and, consequently, compact. As $\beta S$ is a topological inverse semigroup, [37, Proposition 3.8] implies that $\beta S$ is a pospace with respect to the natural partial order. By Proposition 5.1, $\beta S\setminus S$ contains a minimal element $h$. Let $e=hh^{-1}$ and $f=h^{-1}h$. Since $E(S)=E(\beta S)$, we have $e,f\in S$. Let
@@
Fix an open neighborhood $U$ of $h$. Since $ehf=h$, the continuity of the semigroup operation in $\beta S$ yields an open neighborhood $V$ of $h$ such that $eVf\subseteq U$. Observe that for every $x\in S\cap V$ we have $exf\in T\cap U$. Since $U$ was chosen arbitrarily, $h\in\operatorname{cl}_{\beta S}(T)$.
Since $S$ is a topological inverse semigroup, the maximal subgroups $H_{e}=\{x\in S\colon xx^{-1}=e=x^{-1}x\}$ and $H_{f}=\{x\in S\colon xx^{-1}=f=x^{-1}x\}$ are closed in $S$. Theorem 1.3 implies that each closed subgroup of $S$ is compact. Thus, the sets $H_{e}$ and $H_{f}$ are compact. Observe that $H_{e}$ is the $\mathcal{H}$-class of the element $e$ and $H_{f}$ is the $\mathcal{H}$-class of the element $f$. Consider the $\mathcal{H}$-class $H_{e,f}:=\{x\in S\colon xx^{-1}=e\hbox{ and }x^{-1}x=f\}$. If $H_{e,f}\neq\varnothing$, then [36, Proposition 5] implies that $H_{e,f}$ lies in the same $\mathcal{D}$-class with $H_{e}$ and $H_{f}$. By Proposition 5.2, the set $H_{e,f}$ is compact. Thus $h\in\operatorname{cl}_{\beta S}(T\setminus H_{e,f})$. Two cases are possible:
@@
(i) Since $e=hh^{-1}$ and $\beta S$ is a topological inverse semigroup, for each open neighborhood $U$ of $e$ there exists an open neighborhood $V$ of $h$ such that $VV^{-1}\subseteq U$. By the assumption, there is $x\in V$ such that $xx^{-1}<e$ and $xx^{-1}\in U$. It follows that $e\in\operatorname{cl}_{S}({\downarrow}e\setminus\{e\})$. Since $S$ is first-countable, there exists a sequence $\{e_{n}:n\in\omega\}\subseteq{\downarrow}e\setminus\{e\}$ that converges to $e$. Since $eh=h$ we get that the sequence $\{e_{n}h:n\in\omega\}$ converges to $h$. Note that $e_{n}h\neq h$ for each $n\in\omega$, as otherwise $e=hh^{-1}=e_{n}hh^{-1}e_{n}=e_{n}ee_{n}=e_{n}$ for some $n\in\omega$, which contradicts the choice of $e_{n}$. Since $e_{n}hh^{-1}e_{n}h=e_{n}h$, we get that $e_{n}h\leq h$ with respect to the natural partial order on $\beta S$. Since $h$ is a minimal element in $\beta S\setminus S$, we get that $e_{n}h\in S$ for all $n\in\omega$. It follows that $\{e_{n}h:n\in\omega\}$ is an infinite closed discrete subset of $S$, which contradicts the countable compactness of $S$.
(ii) Since $f=h^{-1}h$ and $\beta S$ is a topological inverse semigroup, for each open neighborhood $U$ of $f$ there exists an open neighborhood $V$ of $h$ such that $V^{-1}V\subseteq U$. By the assumption, there is $x\in V$ such that $f>x^{-1}x\in U$. It follows that $f\in\operatorname{cl}_{S}({\downarrow}f\setminus\{f\})$. Since $X$ is first-countable, there exists a sequence $\{f_{n}:n\in\omega\}\subseteq{\downarrow}f\setminus\{f\}$ that converges to $f$. Since $hf=h$ we get that the sequence $\{hf_{n}:n\in\omega\}$ converges to $h$. Similarly as above it can be checked that $h\notin\{hf_{n}:n\in\omega\}$ and $hf_{n}\leq h$ for every $n\in\omega$. Since $h$ is a minimal element in $\beta S\setminus S$, we get that $hf_{n}\in S$ for all $n\in\omega$. It follows that $\{hf_{n}:n\in\omega\}$ is an infinite closed discrete subset of $S$, which contradicts the countable compactness of $S$.
@@
A band is a semigroup consisting of idempotents. A semigroup $S$ is called simple if $S$ contains no proper two-sided ideals or, in other words, $SxS=S$ for any $x\in S$. The following example shows that Theorem A does not generalize over simple bands.
@@
Let $X$ be one of the consistent examples of a locally compact noncompact Nyikos space discussed in the introduction. Let $X_{1}$ be the space $X$ endowed with the left zero operation, i.e., $xy=x$ for each $x,y\in X$. It is easy to see that this operation is continuous and associative, that is, $X_{1}$ is a topological semigroup. Let $X_{2}$ be the space $X$ endowed with the right zero operation, i.e., $xy=y$ for each $x,y\in X$. Similarly, one can check that $X_{2}$ is a topological semigroup. Since the finite product of Nyikos spaces remains Nyikos, the direct product $S:=X_{1}{\times}X_{2}$ is a noncompact Nyikos topological semigroup. Note that $(a,b)(a,b)=(a,b)$ for every $(a,b)\in S$, which makes $S$ a band. Fix $(a,b)\in S$. Then for any $(c,d)\in S$ we have $(c,d)=(c,a)(a,b)(b,d)\in S(a,b)S$, witnessing that $S=S(a,b)S$. Since the point $(a,b)$ was chosen arbitrarily, the band $S$ is simple.
The Ostaszewski space is constructed under ($\diamondsuit$) in [43]. It is a Nyikos space and possesses (among others) the following properties: locally compact, hereditary separable, each compact subspace is countable. The results of this paper establish another peculiar property of the Ostaszewski space.
@@
Let $Y$ be an inverse subsemigroup of $S$. Note that $\overline{Y}$ is a locally compact first-countable countably compact topological subsemigroup of $S$. Since $S$ is hereditary separable, $\overline{Y}$ is a locally compact Nyikos topological semigroup. Corollary 1.4 implies that $\overline{Y}$ is a topological inverse semigroup. By Theorem A, $\overline{Y}$ is compact. Since each compact subspace of the Ostaszewski space is countable, we get $|\overline{Y}|\leq\omega$. Hence $|Y|\leq\omega$, as required. ∎

### v2

@@ 4 Nyikos semilattices
To derive a contradiction assume that $X$ contains a chain $L=\{l_{\alpha}:\alpha<\omega_{1}\}$ such that $l_{\alpha}\leq l_{\beta}$ if and only if $\alpha\leq\beta$. By [23, Corollary 3.10.15], $X{\times}X$ is countable compact and thus pseudocompact. Theorem 2.9 implies that $\beta X$ is a compact topological semigroup. Since $X$ is dense in $\beta X$, we obtain that $\beta X$ is a semilattice. By Lemma 3.4, there exists $z=\sup L\in\beta X$. Lemma 4.7 implies that there exists $l_{\alpha}\in L$ such that for each neighborhood $U$ of $l_{\alpha}$ the set $U\setminus I_{L}$ contains an element $d_{U}$. Since $X$ is locally compact, there exists a neighborhood $W$ of $l_{\alpha}$ such that $\operatorname{cl}_{X}(W)$ is compact. The continuity of the semilattice operation in $\beta X$ yields neighborhoods $U$ of $l_{\alpha}$ and $V$ of $z$ such that $UV\subseteq W$. By Lemma 3.4, there exists an ordinal $\delta<\omega_{1}$ such that $L^{\prime}=\{l_{\xi}:\delta<\xi<\omega_{1}\}\subseteq V$. Note that $d_{U}L^{\prime}\subseteq W$, $L^{\prime}$ is isomorphic to $(\omega_{1},\min)$ and $|d_{U}L^{\prime}|=|d_{U}L|=\omega_{1}$, by the choice of $d_{U}$. Since for each $a\in X$ the map $\phi_{a}:X\rightarrow X$, $\phi_{a}(x)=ax$ is a homomorphism, Lemma 3.7 implies that the chain $d_{U}L^{\prime}$ is isomorphic to $(\beta,\min)$ for some ordinal $\beta\leq\omega_{1}$. Taking into account that $|d_{U}L^{\prime}|=\omega_{1}$, we get that $d_{U}L^{\prime}$ is isomorphic to $(\omega_{1},\min)$. Since $\operatorname{cl}_{X}(W)$ is compact, the chain $\operatorname{cl}_{X}(d_{U}L^{\prime})\subseteq\operatorname{cl}_{X}(W)$ is compact, which contradicts Lemma 3.6. ∎
@@
Seeking a contradiction, assume that a locally compact Nyikos semilattice $X$ contains a chain $L=\{l_{\alpha}:\alpha<\omega_{1}\}$ such that $l_{\alpha}\leq l_{\beta}$ if and only if $\alpha\geq\beta$. By Theorem 2.9, $\beta X$ is a compact topological semigroup. Since $X$ is dense in $\beta X$, we obtain that $\beta X$ is a semilattice. By Lemma 3.5, there exists $z=\inf L\in\beta X\setminus X$. Lemma 4.1 implies that the set $I_{L}=\{x\in X:|xL|<\omega_{1}\}$ is not empty, as it contains the minimum of $X$.
@@
By Lemma 4.5, $I_{L}$ is closed. To derive a contradiction, assume that $I_{L}$ is not open. Then there exists $y\in I_{L}$ such that each neighborhood $U$ of $y$ contains an element $d_{U}\in X\setminus I_{L}$. Let us show that $yz\in X$. Since $y\in I_{L}$, we have $|yL|<\omega_{1}$. Then there exists an uncountable and thus cofinal subset $A\subseteq\omega_{1}$ such that $yl_{\xi}=yl_{\delta}$ for all $\xi,\delta\in A$. Let $\alpha=\min A$. It is routine to check that $yl_{\xi}=yl_{\alpha}$ for each $\xi\geq\alpha$. By Lemma 3.5, for each neighborhood $V$ of $z$ there exists $\delta_{V}<\omega_{1}$ such that $\{l_{\xi}:\xi\geq\delta_{V}\}\subseteq V$. Then $yl_{\alpha}\in yV$ for each neighborhood $V\subseteq\beta X$ of $z$. Since $\beta X$ is a Hausdorff topological semilattice, we get $yz=yl_{\alpha}\in X$. Since $X$ is locally compact, there is a neighborhood $W\subseteq X$ of $yz$ such that $\operatorname{cl}_{X}(W)$ is compact. The continuity of the semilattice operation in $\beta X$ yields neighborhoods $U$ of $y$ and $V$ of $z$ such that $UV\subseteq W$. Let $L^{\prime}=\{l_{\alpha}:\alpha\geq\delta_{V}\}\subseteq V$. Note that $d_{U}L^{\prime}\subseteq W$. By the choice of $d_{U}$, $|d_{U}L|=\omega_{1}$, which implies $|d_{U}L^{\prime}|=\omega_{1}$. As left shifts in $\beta X$ are homomorphisms and $L^{\prime}$ is order isomorphic to $(\omega_{1},\max)$, Lemma 3.7(2) implies that $d_{U}L^{\prime}$ is also order isomorphic to $(\omega_{1},\max)$. Since $\operatorname{cl}_{X}(W)$ is compact, the chain $\operatorname{cl}_{X}(d_{U}L^{\prime})\subseteq\operatorname{cl}_{X}(W)$ is compact, which contradicts Lemma 3.6. ∎
By [23, Corollary 3.6.5], the set $\operatorname{cl}_{\beta X}(I_{L})$ is clopen.
@@
There exists a finite set $F\subseteq X\setminus I_{L}$ such that $\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})=\bigcup_{x\in F}{\uparrow}x$.
@@
Let us first check that $\operatorname{cl}_{\beta X}(I_{L})$ is an ideal in $\beta X$. It suffices to show that $ab\in\operatorname{cl}_{\beta X}(I_{L})$ for any $a\in\operatorname{cl}_{\beta X}(I_{L})$ and $b\in\beta X$. Fix a neighborhood $W$ of $ab$. By the continuity of the semigroup operation in $\beta X$, there exist neighborhoods $U$ of $a$ and $V$ of $b$ such that $UV\subseteq W$. Then there exist $c\in U\cap I_{L}$ and $d\in X\cap V$ such that $cd\in W\cap I_{L}$. Thus $ab\in\operatorname{cl}_{\beta X}(I_{L})$, as required. Consider the Rees quotient semilattice $S=\beta X/{\operatorname{cl}_{\beta X}(I_{L})}$ which is obtained by contracting the clopen ideal $\operatorname{cl}_{\beta X}(I_{L})$ to a point denoted by 0. Obviously, $S$ is a compact topological semilattice, $0=\inf S$ and 0 is isolated in $S$. Further we agree to identify the sets $S\setminus\{0\}$ and $\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})$. Observe that the set $\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})$ is nonempty, as it contains $L$. Let $\mathfrak{C}$ be the set of all maximal chains in $S\setminus\{0\}$. By the compactness of $S\setminus\{0\}$, each $C\in\mathfrak{C}$ contains its minimum. To derive a contradiction, assume that the set
$F=\{\min C:C\in\mathfrak{C}\}\subseteq\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})$
is infinite. It is easy to see that $ab=0$ for any distinct $a,b\in F$. So, $T=F\cup\{0\}$ is a subsemilattice of $S$. Since each chain in $T$ is finite, Theorem 3.2 implies that $T$ is closed in $S$ and thus compact. Consider an accumulation point $y\in T$ of $F$. Observe that for each neighborhood $U$ of $y$ there exist distinct $a,b\in U\cap F$ witnessing that $0=ab\in UU$. Since $S$ is a Hausdorff topological semilattice, we get that $y=yy=0$. On the other hand, $y\neq 0$, as 0 is an isolated point, a contradiction. So the set $F$ is finite. At this point it is straightforward to check that $S\setminus\{0\}=\beta X\setminus\operatorname{cl}_{\beta X}(I_{L})=\bigcup_{x\in F}{\uparrow}x$.
It remains to check that $F\subseteq X$. Fix any $f\in F$. Observe that
@@
is an open subsemilattice of $S$ such that $\inf P=f\in P$. Then $P\cap X$ is an open subsemilattice of $X$. Taking into account that $X$ is separable, Lemma 4.1 yields that the semilattice $\operatorname{cl}_{X}(P\cap X)$ contains the smallest element $p$. Since $\operatorname{cl}_{X}(P\cap X)\subseteq\operatorname{cl}_{\beta X}(P)\subseteq{\uparrow}f$, we get $f\leq p$. Since $X$ is dense in $\beta X$ and $P$ is open, for every neighborhood $U$ of $f$ the set $U\cap P\cap X$ is not empty. Thus $p\in Up$ for each neighborhood $U$ of $f$. Since $S$ is a Hausdorff topological semilattice, $fp=p$, i.e. $p\leq f$. Hence $p=f$ and $f\in\operatorname{cl}_{X}(P\cap X)\subseteq X$. ∎
@@
Since $L\subseteq X\setminus I_{L}$, we obtain
@@
By Claim 4.11, there exists $f\in F\subseteq X\setminus I_{L}$ such that $f\leq z$. But then $fL=(fz)L=f(zL)=\{fz\}=\{f\}$. Hence $f\in I_{L}$, a contradiction. ∎
@@
Propositions 4.8 and 4.9 imply that each locally compact Nyikos topological semilattice $X$ contains no isomorphic copies of $(\omega_{1},\min)$ or $(\omega_{1},\max)$. By Theorem 1.5, $X$ is compact. ∎
5 Proof of the main result and final remarks
@@
For more about Green’s relations on inverse semigroups see [37, Chapter 3.2]. The following two results can be considered folklore.
@@
In fact, Proposition 5.2 follows from the proof of [32, Lemma 2.2.3]. We are in a position to prove Main Theorem 1.4, which states that every locally compact Nyikos inverse topological semigroup is compact.
Proof of Main Theorem 1.4.
Consider a locally compact Nyikos inverse topological semigroup $S$. By Corollary 1.7, $S$ is a topological inverse semigroup. Note that the semilattice of idempotents $E(S)$ is a retract of $S$ under the continuous map $\pi:x\mapsto xx^{-1}$. It follows that $E(S)$ is a locally compact Nyikos topological semilattice. Theorem 4.12 implies that $E(S)$ is compact. By Theorem 1.6, $\beta S$ is a topological inverse semigroup. We claim that $E(S)=E(\beta S)$. Since $E(S)$ is compact, it suffices to show that $E(S)$ is dense in $E(\beta S)$. For this fix any $e\in E(\beta S)$ and its neighborhood $U$. Since $ee^{-1}=e$ and $\beta S$ is a topological inverse semigroup, there exists a neighborhood $V$ of $e$ such that $VV^{-1}\subseteq U$. Since $S$ is dense in $\beta S$, there exists $s\in S\cap V$. It follows that $ss^{-1}\in U\cap E(S)$. Since the neighborhood $U$ is chosen arbitrarily, we get that $e\in\operatorname{cl}_{\beta S}(E(S))$. Hence $E(S)$ is dense in $E(\beta S)$, which implies $E(S)=E(\beta S)$, as claimed.
Seeking a contradiction, assume that $\beta S\setminus S\neq\varnothing$. Since $S$ is locally compact, the remainder $\beta S\setminus S$ is closed and, consequently, compact. As $\beta S$ is a topological inverse semigroup, [38, Proposition 3.8] implies that $\beta S$ is a pospace with respect to the natural partial order. By Proposition 5.1, $\beta S\setminus S$ contains a minimal element $h$. Let $e=hh^{-1}$ and $f=h^{-1}h$. Since $E(S)=E(\beta S)$, we have $e,f\in S$. Let
@@
Fix a neighborhood $U$ of $h$. Since $ehf=h$, the continuity of the semigroup operation in $\beta S$ yields a neighborhood $V$ of $h$ such that $eVf\subseteq U$. Observe that for every $x\in S\cap V$ we have $exf\in T\cap U$. Since $U$ was chosen arbitrarily, $h\in\operatorname{cl}_{\beta S}(T)$.
Since $S$ is a topological inverse semigroup, the subgroups $H_{e}=\{x\in S\colon xx^{-1}=e=x^{-1}x\}$ and $H_{f}=\{x\in S\colon xx^{-1}=f=x^{-1}x\}$ are closed in $S$. Theorem 1.3 implies that each closed subgroup of $S$ is compact. Thus, the sets $H_{e}$ and $H_{f}$ are compact. Observe that $H_{e}$ is the $\mathcal{H}$-class of the element $e$ and $H_{f}$ is the $\mathcal{H}$-class of the element $f$. Consider the $\mathcal{H}$-class $H_{e,f}:=\{x\in S\colon xx^{-1}=e\hbox{ and }x^{-1}x=f\}$. If $H_{e,f}\neq\varnothing$, then [37, Proposition 5] implies that $H_{e,f}$ lies in the same $\mathcal{D}$-class with $H_{e}$ and $H_{f}$. By Proposition 5.2, the set $H_{e,f}$ is compact. Thus $h\in\operatorname{cl}_{\beta S}(T\setminus H_{e,f})$. Two cases are possible:
@@
(i) Since $e=hh^{-1}$ and $\beta S$ is a topological inverse semigroup, for each neighborhood $U$ of $e$ there exists a neighborhood $V$ of $h$ such that $VV^{-1}\subseteq U$. By assumption, there is $x\in V$ such that $xx^{-1}<e$ and $xx^{-1}\in U$. It follows that $e\in\operatorname{cl}_{S}({\downarrow}e\setminus\{e\})$. Since $S$ is first-countable, there exists a sequence $\{e_{n}:n\in\omega\}\subseteq{\downarrow}e\setminus\{e\}$ that converges to $e$. Since $eh=h$ we get that the sequence $\{e_{n}h:n\in\omega\}$ converges to $h$. Note that $e_{n}h\neq h$ for each $n\in\omega$, as otherwise $e=hh^{-1}=e_{n}hh^{-1}e_{n}=e_{n}ee_{n}=e_{n}$ for some $n\in\omega$, which contradicts the choice of $e_{n}$. Since $e_{n}hh^{-1}e_{n}h=e_{n}h$, we get that $e_{n}h\leq h$ with respect to the natural partial order on $\beta S$. Since $h$ is a minimal element in $\beta S\setminus S$, we get that $e_{n}h\in S$ for all $n\in\omega$. It follows that $\{e_{n}h:n\in\omega\}$ is an infinite closed discrete subset of $S$, which contradicts the countable compactness of $S$.
(ii) Since $f=h^{-1}h$ and $\beta S$ is a topological inverse semigroup, for each neighborhood $U$ of $f$ there exists a neighborhood $V$ of $h$ such that $V^{-1}V\subseteq U$. By assumption, there is $x\in V$ such that $f>x^{-1}x\in U$. It follows that $f\in\operatorname{cl}_{S}({\downarrow}f\setminus\{f\})$. Since $X$ is first-countable, there exists a sequence $\{f_{n}:n\in\omega\}\subseteq{\downarrow}f\setminus\{f\}$ that converges to $f$. Since $hf=h$ we get that the sequence $\{hf_{n}:n\in\omega\}$ converges to $h$. Similarly as above it can be checked that $h\notin\{hf_{n}:n\in\omega\}$ and $hf_{n}\leq h$ for every $n\in\omega$. Since $h$ is a minimal element in $\beta S\setminus S$, we get that $hf_{n}\in S$ for all $n\in\omega$. It follows that $\{hf_{n}:n\in\omega\}$ is an infinite closed discrete subset of $S$, which contradicts the countable compactness of $S$.
@@
A band is a semigroup consisting of idempotents. A semigroup $S$ is called simple if $S$ contains no proper two-sided ideals or, in other words, $SxS=S$ for any $x\in S$. The following example shows that Main Theorem 1.4 does not generalize over simple bands.
@@
Let $X$ be one of the consistent examples of a locally compact noncompact Nyikos space discussed in the introduction. Let $X_{1}$ be the space $X$ endowed with the left zero operation, i.e., $xy=x$ for each $x,y\in X$. It is easy to see that this operation is continuous and associative, that is, $X_{1}$ is a topological semigroup. Let $X_{2}$ be the space $X$ endowed with the right zero operation, i.e., $xy=y$ for each $x,y\in X$. Similarly, one can check that $X_{2}$ is a topological semigroup. Taking into account [23, Corollary 3.10.15], it is easy to check that the finite product of Nyikos spaces remains Nyikos. Hence the direct product $S:=X_{1}{\times}X_{2}$ is a noncompact Nyikos topological semigroup. Note that $(a,b)(a,b)=(a,b)$ for every $(a,b)\in S$, which makes $S$ a band. Fix $(a,b)\in S$. Then for any $(c,d)\in S$ we have $(c,d)=(c,a)(a,b)(b,d)\in S(a,b)S$, witnessing that $S=S(a,b)S$. Since the point $(a,b)$ was chosen arbitrarily, the band $S$ is simple.
The Ostaszewski space was constructed under ($\diamondsuit$) in [44]. It is a Nyikos space and possesses (among others) the following properties: locally compact, hereditary separable, each compact subspace is countable. The results of this paper establish another peculiar property of the Ostaszewski space.
@@
Let $Y$ be an inverse subsemigroup of $S$. Note that $\overline{Y}$ is a locally compact first-countable countably compact topological subsemigroup of $S$. Since $S$ is hereditary separable, $\overline{Y}$ is a locally compact Nyikos topological semigroup. Corollary 1.7 implies that $\overline{Y}$ is a topological inverse semigroup. By Main Theorem 1.4, $\overline{Y}$ is compact. Since each compact subspace of the Ostaszewski space is countable, we get $|\overline{Y}|\leq\omega$. Hence $|Y|\leq\omega$, as required. ∎